# Dark Deceptions in DHCP: Dismantling Network Defenses

### v1

@@ Abstract
DHCP Vulnerabilities, Network Security, VPN Exploits, Positive Unlabeled Learning
@@ II-B The CIA Triad
The CIA triad encapsulates the core principles of cybersecurity: Confidentiality, Integrity, and Availability. DHCP attacks often undermine these principles, necessitating robust controls to safeguard network assets.
@@ V-C Implications and Countermeasures
Addressing this requires a multi-layered approach. Users must adopt VPN solutions that rigorously enforce route integrity, including safeguards against arbitrary route injection. System administrators can mitigate such risks by implementing DHCP snooping and using MAC binding to prevent unauthorized DHCP servers from operating within the network. Additionally, promoting best practices such as routinely updating VPN software, enabling necessary protocols like kill-switches, and enabling DNS leak protection can close potential attack vectors.

### v2

@@ Abstract
DHCP Vulnerabilities, Network Security, VPN Exploits, TunnelVision, Positive Unlabeled Learning
@@ II-B The CIA Triad
The CIA Triad encapsulates the core principles of cybersecurity: Confidentiality, Integrity, and Availability.
  • Confidentiality: Ensuring information is accessible only to authorized users.
  • Integrity: Maintaining the accuracy and trustworthiness of data.
  • Availability: Guaranteeing information is accessible when required.
The three types of controls–or measures used to manage risk–that can be applied to address the above principles are:
  • Corrective: Actions taken to restore systems after an incident.
  • Detective: Measures to identify and alert on security issues.
  • Mitigative: Strategies to reduce the impact of potential threats.
DHCP attacks often undermine these principles, necessitating robust controls to safeguard network assets.
@@ V-C Implications and Countermeasures
Addressing this requires a multi-layered approach. Users must adopt VPN solutions that rigorously enforce route integrity, including safeguards against arbitrary route injection. System administrators can mitigate such risks by implementing DHCP snooping and using MAC binding to prevent unauthorized DHCP servers from operating within the network. Additionally, promoting best practices such as routinely updating VPN software, enabling necessary protocols like kill-switches [6, 7, 8], and enabling DNS leak protection can close potential attack vectors.